# A Study on Zero-shot Non-intrusive Speech Assessment using Large Language Models

### v1

@@ Abstract
This work investigates two strategies for zero-shot non-intrusive speech assessment leveraging large language models. First, we explore the audio analysis capabilities of GPT-4o. Second, we propose GPT-Whisper, which uses Whisper as an audio-to-text module and evaluates the text’s naturalness via targeted prompt engineering. We evaluate assessment metrics predicted by GPT-4o and GPT-Whisper examining their correlations with human-based quality and intelligibility assessments, and character error rate (CER) of automatic speech recognition. Experimental results show that GPT-4o alone is not effective for audio analysis; whereas, GPT-Whisper demonstrates higher prediction, showing moderate correlation with speech quality and intelligibility, and high correlation with CER. Compared to supervised non-intrusive neural speech assessment models, namely MOS-SSL and MTI-Net, GPT-Whisper yields a notably higher Spearman’s rank correlation with Whisper’s CER. These findings validate GPT-Whisper as a reliable method for accurate zero-shot speech assessment without requiring additional training data (speech data and corresponding assessment scores).
@@ I Introduction
Speech assessment metrics play a critical role in evaluating a variety of speech-related applications, including speech enhancement [1, 2], hearing aid (HA) devices [3, 4, 5], and telecommunications [6]. With the advancement of deep learning and the need for accurate non-intrusive speech assessment metrics, researchers have increasingly adopted deep learning models for speech assessment [7, 8, 9, 10, 11, 12, 13]. To achieve accurate automatic assessments, various strategies have been explored, such as reducing listener bias [14], integrating large pre-trained models (e.g., self-supervised learning (SSL) models [15, 16], Whisper [11], and speech language models[17]), utilizing ensemble learning [7, 8, 9], and incorporating pseudo labels [18]. Despite significant performance improvements, achieving satisfactory generalization with limited training samples remains challenging.
@@
In this paper, we explore two strategies for zero-shot non-intrusive speech assessment by leveraging large language models (LLM). First, we directly leverage the audio analysis capabilities of GPT-4o for speech assessment. Second, we propose a more advanced approach, namely GPT-Whisper. Specifically, we use Whisper [22] as an audio-to-text module to create text representations, which are then assessed using targeted prompt engineering focused on evaluating the naturalness of the predicted text. To the best of our knowledge, this is the first attempt to leverage ChatGPT for speech assessment. Furthermore, to gain a deeper insight into our new assessment metric computed by GPT-Whisper, we also evaluate its correlation with several metrics, including human-based quality assessment, human-based intelligibility assessment, and the character error rate (CER) of automatic speech recognition (ASR) models. We use two open-source ASR models: Whisper [22], and Google ASR [23] to generate CER scores. Finally, we compare our approach with two supervised neural speech assessment models: MOS-SSL [15] and MTI-Net [24].
Our experimental results first confirm that using ChatGPT alone for audio analysis is not optimal because it mainly relies on factors such as amplitude range, standard deviation, and signal-to-noise ratio (SNR) to assess the quality or intelligibility of audio data. Next, in terms of Spearman’s rank correlation coefficient (SRCC) [25], our GPT-Whisper metric yields moderate correlation with human-based quality assessment (0.4360), human-based intelligibility assessment (0.5485), and Google ASR’s CER (0.5415) and high correlation with Whisper’s CER (0.7784). This further validates our new metric as a viable assessment for evaluating speech in a zero-shot manner, especially given its higher correlation with CER. Finally, when comparing GPT-Whisper with MOS-SSL and all variants of MTI-Net (wav2vec and Whisper) in predicting Whisper’s CER, experimental results show that GPT-Whisper outperforms them in terms of SRCC (0.7784 vs. 0.7482, 0.7418, and 0.7655).
@@ III-A Experimental setup
Two evaluation metrics, linear correlation coefficient (LCC) and Spearman’s rank correlation coefficient (SRCC) [25], are used to measure the performance of the proposed methods. Higher LCC and SRCC values indicate a stronger correlation between the predicted score and the ground-truth score, indicating better assessment prediction performance.
@@ III-C1 Correlation analysis
In initial experiments, we estimate GPT-Whisper scores using the approach outlined in Section II. Experimental results show that GPT-4o exhibits notable reasoning capabilities in text comprehension and demonstrates solid contextual understanding, as shown in Fig. 4. GPT-4o effectively evaluates the coherence of text inputs and determines whether connections are natural or unnatural. Based on these results, we conclude that the scores produced by GPT-Whisper are more reliable than those produced previously by GPT-4o for speech assessment.
Based on the promising performance of GPT-Whisper, we intend to further investigate which assessment metrics are more correlated with the GPT-Whisper score. We assume that GPT-Whisper can be used as an alternative in scenarios where there is no corresponding label, and the training data is insufficient to establish a deep learning-based speech assessment model. In our evaluation, we compare GPT-Whisper scores with human-based quality, human-based intelligibility scores, and the CER of two ASR models (Whisper and Google ASR). The results in Fig. 5 show that the GPT-Whisper metric has a higher correlation with intelligibility-based metrics than the quality one and has the highest correlation with the CER of Whisper ASR. These results further confirm GPT-Whisper’s ability to perform zero-shot speech assessment and serve as a surrogate metric for evaluating speech recognition performance, given its high correlation score with the CER of Whisper (SRCC of 0.7784).
@@
Figure 5: Correlation of GPT-Whisper scores with other metrics.
@@ III-C2 Comparison with supervised methods
We now compare GPT-Whisper with two supervised speech assessment models, namely MOS-SSL [15], and MTI-Net [24], on the task of predicting the CER of Whisper. We use Whisper’s CER as the ground truth because the GPT-Whisper score has a higher correlation with this CER score than with human-based quality and intelligibility. Therefore, it is interesting to compare GPT-Whisper with these supervised models on this metric. Specifically, MOS-SSL incorporates wav2vec 2.0 to predict MOS scores, while MTI-Net employs cross-domain features (raw waveform, spectral features, and SSL features) to predict intelligibility metrics (intelligibility, CER, and STOI) using multi-task learning [24]. We trained MOS-SSL and MTI-Net using the TMHINT-QI(s) training set. MOS-SSL was trained to predict Whisper’s CER, while MTI-Net was trained to predict intelligibility, Google ASR’s CER, Whisper’s CER, and STOI. Given the significant efficacy of integrating Whisper representations into robust speech assessment models, [11, 8, 9], we also integrated Whisper into the MTI-Net model. Therefore, we prepared two versions: MTI-Net (wav2vec) and MTI-Net (Whisper).
The results in Table I confirm that GPT-Whisper performs comparably to supervised non-intrusive speech assessment models. Interestingly, GPT-Whisper outperforms MOS-SSL and MTI-Net (wav2vec) in LCC and SRCC. Additionally, GPT-Whisper has higher SRCC than MTI-Net (Whisper), although its LCC is slightly lower. These results demonstrate the effectiveness of GPT-Whisper as a zero-shot speech assessment model, showing comparable or even superior performance to supervised methods.
TABLE I: Comparison of GPT-Whisper with supervised methods.
@@ III-C3 Score distribution analysis
In this section, we further analyze the distribution of scores evaluated by GPT-Whisper and MTI-Net (Whisper) and the corresponding CER labels of Whisper. This analysis aims to evaluate whether GPT-Whisper provides accurate score estimations across the entire score range. As shown in Fig. 6, GPT-Whisper successfully predicts the entire range of scores and shows a more similar distribution to that of CER generated by the Whisper ASR, especially for CER less than 0.25. The closeness of the distribution of predicted CER scores to the distribution of true CER scores validates GPT-Whisper as a new assessment metric and the potential of using ChatGPT for speech assessment by implementing the GPT-Whisper method.
@@ IV Conclusions
In this paper, we have proposed two strategies for zero-shot non-intrusive speech assessment. The first strategy leverages the audio analysis capabilities of GPT-4o. The second strategy introduces a more advanced approach that uses Whisper as an audio-to-text module and evaluates the naturalness of the generated text through targeted prompt engineering, so the model is called GPT-Whisper. To the best of our knowledge, this is the first attempt to leverage ChatGPT for speech assessment. Experimental results confirm that GPT-4o alone is insufficient for accurate speech assessment. In contrast, GPT-Whisper demonstrates potential as a zero-shot speech assessment method, with yielding moderate correlation with human-based quality assessment (SRCC of 0.4360), human-based intelligibility assessment (SRCC of 0.5485), and Google ASR’s CER (SRCC of 0.5415) and notably high correlation with Whisper’s CER (SRCC of 0.7784). Furthermore, in predicting Whisper’s CER, GPT-Whisper outperforms the supervised models MOS-SSL, MTI-Net (wav2vec), and MTI-Net (Whisper) in terms of SRCC (0.7784 vs. 0.7482, 0.7418, and 0.7655). This confirms the ability of GPT-Whisper as a reliable zero-shot speech assessment method and the potential of using ChatGPT for speech assessment through the GPT-Whisper method. In the future, we plan to explore more advanced prompt engineering to further leverage ChatGPT to estimate other speech assessment metrics. We also intend to investigate direct integration of GPT-Whisper with speech-processing applications to enhance their performance.

### v2

@@ Abstract
This work investigates two strategies for zero-shot non-intrusive speech assessment leveraging large language models. First, we explore the audio analysis capabilities of GPT-4o. Second, we propose GPT-Whisper, which uses Whisper as an audio-to-text module and evaluates the text’s naturalness via targeted prompt engineering. We evaluate the assessment metrics predicted by GPT-4o and GPT-Whisper, examining their correlation with human-based quality and intelligibility assessments and the character error rate (CER) of automatic speech recognition. Experimental results show that GPT-4o alone is less effective for audio analysis, while GPT-Whisper achieves higher prediction accuracy, has moderate correlation with speech quality and intelligibility, and has higher correlation with CER. Compared to SpeechLMScore and DNSMOS, GPT-Whisper excels in intelligibility metrics, but performs slightly worse than SpeechLMScore in quality estimation. Furthermore, GPT-Whisper outperforms supervised non-intrusive models MOS-SSL and MTI-Net in Spearman’s rank correlation for Whisper’s CER. These findings validate GPT-Whisper’s potential for zero-shot speech assessment without requiring additional training data.
@@ I Introduction
Speech assessment metrics play a critical role in evaluating a variety of speech-related applications, including speech enhancement [1, 2], hearing aid (HA) devices [3, 4, 5], and telecommunications [6]. With the advancement of deep learning and the need for accurate non-intrusive speech assessment metrics, researchers have increasingly adopted deep learning models for speech assessment [7, 8, 9, 10, 11, 12, 13, 14, 15]. To achieve accurate automatic assessments, various strategies have been explored, such as reducing listener bias [16], integrating large pre-trained models (e.g., self-supervised learning (SSL) models [17, 18], Whisper [11], and speech language models[19]), utilizing ensemble learning [7, 8, 9], and incorporating pseudo labels [20]. Despite significant performance improvements, achieving satisfactory generalization with limited training samples remains challenging.
@@
In this paper, we explore two strategies for zero-shot non-intrusive speech assessment by leveraging large language models (LLM). First, we directly leverage the audio analysis capabilities of GPT-4o for speech assessment. Second, we propose a more advanced approach, namely GPT-Whisper. Specifically, we use Whisper [24] as an audio-to-text module to create text representations, which are then assessed using targeted prompt engineering focused on evaluating the naturalness of the predicted text. To gain a deeper insight into our new assessment metric computed by GPT-Whisper, we also evaluate its correlation with several metrics, including human-based quality assessment, human-based intelligibility assessment, and the character error rate (CER) of automatic speech recognition (ASR) models. We use two open-source ASR models: Whisper [24] and Google ASR [25] to generate CER scores. To the best of our knowledge, this is the first attempt to leverage ChatGPT to simultaneously estimate speech quality, intelligibility, and CER. Finally, we compare our approach with DNSMOS [26], one unsupervised neural speech assessment model—SpeechLMscore [19]—and two supervised neural speech assessment models—MOS-SSL [17] and MTI-Net [27].
Our experimental results first confirm that using ChatGPT alone for audio analysis is less optimal because it mainly relies on factors such as amplitude range, standard deviation, and signal-to-noise ratio (SNR) to assess the quality or intelligibility of audio data. Next, in terms of Spearman’s rank correlation coefficient (SRCC) [28], our GPT-Whisper metric yields moderate correlation with human-based quality assessment (0.4360), human-based intelligibility assessment (0.5485), and Google ASR’s CER (0.5415) and high correlation with Whisper’s CER (0.7784). This further validates our metric’s capability for zero-shot evaluation, especially given its higher correlation with CER. Next, when comparing GPT-Whisper with DNSMOS, and SpeechLMScore, GPT-Whisper excels in intelligibility metrics (0.5485 vs. 0.2643, 0.0192) while performing slightly below SpeechLMScore in quality estimation (0.5108 vs. 0.4360). Moreover, GPT-Whisper surpasses supervised non-intrusive models, including MOS-SSL and all variants of MTI-Net (wav2vec and Whisper), in predicting Whisper’s CER. Experimental results show that GPT-Whisper achieves a higher SRCC (0.7784) compared to 0.7482, 0.7418, and 0.7655 for the other models.
@@ III-A Experimental setup
Two evaluation metrics, namely linear correlation coefficient (LCC) and Spearman’s rank correlation coefficient (SRCC) [28], are used to measure the performance of the proposed methods. Higher LCC and SRCC values indicate a stronger correlation between the predicted score and the ground-truth score, indicating better assessment prediction performance.
@@ III-C1 Correlation analysis
In initial experiments, we estimate GPT-Whisper scores using the approach outlined in Section II. Experimental results show that GPT-4o exhibits notable reasoning capabilities in text comprehension and demonstrates solid contextual understanding, as shown in Fig. 4. GPT-4o effectively evaluates the coherence of text inputs and determines whether connections are natural or unnatural.
Given the above results, we aim to further investigate which assessment metrics are more correlated with the GPT-Whisper score. In our evaluation, we compare GPT-Whisper scores with human-based quality and intelligibility scores, as well as the CER of two ASR models (Whisper and Google ASR). We also deployed another variant of GPT-Whisper, namely GPT-Google, where a different ASR module is used. The results in Table I confirm that the GPT-Whisper metric achieves better overall performance than GPT-Google across nearly all metrics, emphasizing the critical role of ASR module robustness in accurate estimation. Furthermore, GPT-Whisper shows a higher correlation with intelligibility-based metrics than with quality metrics, and it has the higher correlation with the CER of Whisper ASR. Finally, when compared with other models such as DNSMOS [26] and SpeechLMScore [19], GPT-Whisper excels in intelligibility metrics, performs slightly below SpeechLMScore in quality estimation, and outperforms DNSMOS across all evaluation metrics.
@@
TABLE I: Comparison GPT-Whisper with other methods for estimating speech quality, speech intelligibility, Google’s CER, and Whisper’s CER.
Model Speech Quality Speech Intelligibility
LCC SRCC LCC SRCC
GPT-Whisper 0.4303 0.4360 0.5226 0.5485
GPT-Google (ours) 0.4188 0.4180 0.4641 0.4857
DNSMOS [26] 0.3088 0.2922 0.0066 0.0192
SpeechLMScore [19] 0.5480 0.5108 0.2741 0.2643
CER Google CER Whisper
GPT-Whisper 0.5469 0.5414 0.7541 0.7784
GPT-Google (ours) 0.6277 0.6962 0.4635 0.4778
SpeechLMScore [19] 0.2460 0.2070 0.1520 0.1126
@@ III-C2 Comparison with supervised methods
We compare GPT-Whisper with two supervised speech assessment models, MOS-SSL [17] and MTI-Net [27], for predicting Whisper’s CER. Specifically, Whisper’s CER is used as the ground truth to evaluate the predictive performance of GPT-Whisper and the comparative assessment models. For the two comparative models, MOS-SSL incorporates wav2vec 2.0 to predict MOS scores, while MTI-Net employs cross-domain features (raw waveform, spectral features, and SSL features) to predict intelligibility metrics (intelligibility, CER, and STOI) using multi-task learning [27]. We trained MOS-SSL and MTI-Net using the TMHINT-QI(s) training set. MOS-SSL was trained to predict Whisper’s CER, while MTI-Net was trained to predict intelligibility, Google ASR’s CER, Whisper’s CER, and STOI. Given the significant efficacy of integrating Whisper representations into robust speech assessment models [11, 8, 9], we also integrated Whisper into the MTI-Net model. Therefore, we prepared two versions: MTI-Net (wav2vec) and MTI-Net (Whisper).
The results in Table II confirm that GPT-Whisper performs comparably to supervised non-intrusive speech assessment models. Interestingly, GPT-Whisper outperforms MOS-SSL and MTI-Net (wav2vec) in LCC and SRCC. Additionally, GPT-Whisper has higher SRCC than MTI-Net (Whisper), although its LCC is slightly lower. These results demonstrate the capability of GPT-Whisper as a zero-shot speech assessment model, showing comparable or even superior performance to supervised methods.
TABLE II: Comparison of GPT-Whisper with supervised methods for estimating Whisper’s CER.
@@ III-C3 Score distribution analysis
In this section, we further analyze the distribution of scores evaluated by GPT-Whisper and MTI-Net (Whisper) and the corresponding CER labels of Whisper. This analysis aims to evaluate whether GPT-Whisper provides accurate score estimations across the entire score range. As shown in Fig. 5, GPT-Whisper successfully predicts the entire range of scores and shows a more similar distribution to that of CER generated by the Whisper ASR, especially for CER less than 0.25. The closeness of the predicted CER score distribution to the true CER score distribution validates GPT-Whisper as a new assessment metric and demonstrates the potential of using a large language model for speech assessment through the GPT-Whisper method.
@@ IV Conclusions
In this paper, we have proposed two strategies for zero-shot non-intrusive speech assessment. The first strategy leverages the audio analysis capabilities of GPT-4o. The second strategy introduces a more advanced approach that uses Whisper as an audio-to-text module and evaluates the naturalness of the generated text through targeted prompt engineering, so the model is called GPT-Whisper. To the best of our knowledge, this is the first attempt to leverage ChatGPT to simultaneously estimate speech quality, intelligibility, and CER. Experimental results confirm that GPT-4o alone is insufficient for accurate speech assessment. In contrast, GPT-Whisper demonstrates potential as a zero-shot speech assessment method, with yielding moderate correlation with human-based quality assessment (SRCC of 0.4360), human-based intelligibility assessment (SRCC of 0.5485), and Google ASR’s CER (SRCC of 0.5415) and notably high correlation with Whisper’s CER (SRCC of 0.7784). Furthermore, compared to SpeechLMScore and DNSMOS, GPT-Whisper excels in intelligibility metrics and performs slightly below SpeechLMScore in quality estimation. Finally, in predicting Whisper’s CER, GPT-Whisper outperforms the supervised models MOS-SSL, MTI-Net (wav2vec), and MTI-Net (Whisper) in terms of SRCC (0.7784 vs. 0.7482, 0.7418, and 0.7655). This confirms GPT-Whisper’s capability as a zero-shot speech assessment method and highlights the potential of large language models for advancing speech evaluation methods. In the future, we plan to explore more advanced prompt engineering to further leverage large language models to estimate other speech assessment metrics. We also intend to investigate direct integration of GPT-Whisper with speech-processing applications to enhance their performance.